OpenStudy (anonymous):

Solve -6(5-2z) + 5z = 8z - 3(16-3z).Please show all your work?

OpenStudy (anonymous):

-6(5-2z) + 5z = 8z - 3(16-3z) -30 + 12z + 5z = 8z -48 +6z 17z= - 18 + 14z 3z = -18 z= -6